the size of a brick is 25 cm by 12 cm by 55 cm and its density is the 1.8 by 10^3 kg/m^3.how many trips must a 3 ton dumper carry out to transport 15000 bricks to a constrution site?

Answer by mananth(16946)

Volume of brick = 0.25m * 0.12m * 0.55m = 0.0165 cu.m
volume of 1 brick = 0.0165 cu.m
density = 1.8*10^3 kg/m^3
weight of 1 brick = 0.0165* 1.8*10^3
29.7 kg
15000 bricks will weigh 15000*29.7
=445500 kg
= 445.5 tons
number of trips = 445.5/3 = 1485 trips